Seafoam:

Jude says that the volume of a square pyramid with base edges of 9.7 in and a height of 9 in is equal to the volume of a cylinder with a radius of 5.47 in and a height of 3 in. Jude rounded his answers to the nearest whole number. Examine Jude's calculations. Is he correct? Volume of Square Pyramid Volume of Cylinder V = one third B(h) V = πr2h V = one third(94.09)(9) V = π(5.472)(3) V = one third(846.81) V = π(29.9209)(3) V = 282 in3 V = π(89.7627) V ≈ 282 in3 Yes, his calculations are correct and the volumes for figures are equal. No, he made a mistake in solving for the volume of the cylinder. Yes, but he made a mistake in solving for the volume of the square pyramid. No, he made a mistake in solving for the volume of both figures.

Seafoam:

No, he made a mistake in solving for the volume of the cylinder.

AZ:

Mhmm, I don't think so Where do you think the mistake is?

AZ:

Volume of Square Pyramid V = one third B(h) B = 9.7 * 9.7 h = 9 multiply all the numbers and look at each step, does it look good?

AZ:

Volume of Cylinder V = πr^2h r = 5.47 h = 3 multiply all the numbers and look at each step, does it look good?

Seafoam:

282.27?

Seafoam:

Oh there equal? thats interesting

AZ:

yeah, there was no mistakes

Seafoam:

Ok great thank you!

AZ:

No problem!